OpenStudy (anonymous):

calculate the derivative of f(x)= (7x-11)+6[(x^.25)+13] is the answer: 7+1.5x?

OpenStudy (anonymous):

7 is ok, 2nd part must be 6*.25(x)^(-0.75)

OpenStudy (anonymous):

mik.wind is right

OpenStudy (anonymous):

can you explain why?

OpenStudy (anonymous):

(7 x - 11) + 6 (x^(25/100) + 13) The above simplified is:\[67+6 x^{1/4}+7 x \]The derivative of the above with respect to x is:\[7+\frac{3}{2 x^{3/4}} \]\[\int\limits \left(7+\frac{3}{2 x^{3/4}}\right) \, dx=6 x^{1/4}+7 x+c \]